# Generative AI Priors Increase Effective Sample Size in Clinical Trials

### v1

@@ 1 Introduction
Our empirical evaluation uses real clinical trial data from NCT00617669, a multi-center non-small cell lung cancer (NSCLC) study, providing a realistic testbed for methodology validation. Notably, this dataset was also used in Barmaz and Ménard (2021), allowing for direct comparison with established meta-analytical priors. Through systematic temperature sensitivity analysis and rigorous cross-validation, we demonstrate the practical utility of LLM-informed priors for clinical safety modeling, benchmarking their performance against the widely used meta-analytical prior framework. Crucially, our results reveal that LLM-based priors can enable substantial reductions in required sample size, demonstrating the potential of LLMs to reduce the number of patients required for statistically robust clinical trials.
@@ 2.2 LLM-Based Prior Elicitation
    2- Disease: Non-small cell lung cancer (NSCLC)
@@
  • All instructions regarding data sources and prior information refer specifically to “NSCLC control arms in multi-center oncology trials” rather than general clinical trial data.
@@ 3.1 Dataset
Our evaluation uses real individual patient data from the control arm of the multi-center clinical trial NCT00617669 (non-small cell lung cancer, NSCLC), which is publicly available via the Project Data Sphere platform. The dataset was originally curated and described by Barmaz and Ménard (2021), and includes adverse event counts for 468 patients across 125 clinical sites, providing a realistic testbed for hierarchical AE modeling. Patient-level AE counts range from 0 to 140, with substantial heterogeneity across sites (mean site size: 3.74 patients, range: 1-27). Notably, this dataset is identical to that used in Barmaz and Ménard (2021), enabling direct methodological comparison to established work.
@@ 5 Discussion
Limitations of this study include the focus on a single disease area (NSCLC) and dataset, as well as the use of only two LLM architectures. Future research should explore validation across diverse clinical contexts, comparison of additional LLMs, and formal assessment of prior informativeness and clinical impact. Integration with regulatory frameworks and extension to other types of clinical endpoints are also important directions for further work.
@@ 6 Conclusion and Future Work
This study demonstrates that LLM-informed prior elicitation, using models such as Llama 3.3 and MedGemma, can substantially improve predictive performance and sample efficiency in hierarchical Bayesian modeling for clinical trials. Through rigorous cross-validation and sample efficiency experiments on real NSCLC trial data, we showed that LLM-derived priors consistently outperform traditional meta-analytical baselines, even when the available training data is limited. Llama 3.3 (Blind, $T=1.0$) provided the best overall results, and both LLMs enabled robust inference with fewer patients, highlighting the practical value of generative AI for clinical safety analysis.
@@ Appendix E Full Disease-Informed Prompt
3TASK: Provide ONLY rate parameters for exponential priors based on NSCLC control arm data.
@@
6- Disease: Non-small cell lung cancer (NSCLC)
@@
17- Use your expert knowledge and draw on published clinical trials, empirical data, or established domain knowledge specific to NSCLC control arms to set informative (not weakly-informative or non-informative) prior rates.
18- Avoid using vague or default values. Base your answer on realistic NSCLC control arm data or strong prior experience relevant to typical AE rates in multi-center oncology trials.

### v2

# How many patients could we save with LLM priors?

@@ 1 Introduction
Our empirical evaluation uses real clinical trial data from NCT00617669, a multi-center metastatic hormone-resistant prostate cancer (HRPC) study, providing a realistic testbed for methodology validation. Notably, this dataset was also used in Barmaz and Ménard (2021), allowing for direct comparison with established meta-analytical priors. Through systematic temperature sensitivity analysis and rigorous cross-validation, we demonstrate the practical utility of LLM-informed priors for clinical safety modeling, benchmarking their performance against the widely used meta-analytical prior framework. Crucially, our results reveal that LLM-based priors can enable substantial reductions in required sample size, demonstrating the potential of LLMs to reduce the number of patients required for statistically robust clinical trials.
@@ 2.2 LLM-Based Prior Elicitation
    2- Disease: Metastatic hormone-resistant prostate cancer (HRPC)
@@
  • All instructions regarding data sources and prior information refer specifically to “HRPC control arms in multi-center oncology trials” rather than general clinical trial data.
@@ 3.1 Dataset
Our evaluation uses real individual patient data from the control arm of the multi-center clinical trial NCT00617669 (metastatic hormone-resistant prostate cancer, HRPC), which is publicly available via the Project Data Sphere platform. The dataset was originally curated and described by Barmaz and Ménard (2021), and includes adverse event counts for 468 patients across 125 clinical sites, providing a realistic testbed for hierarchical AE modeling. Patient-level AE counts range from 0 to 140, with substantial heterogeneity across sites (mean site size: 3.74 patients, range: 1-27). Notably, this dataset is identical to that used in Barmaz and Ménard (2021), enabling direct methodological comparison to established work.
@@ 5 Discussion
Limitations of this study include the focus on a single disease area (HRPC) and dataset, as well as the use of only two LLM architectures. Future research should explore validation across diverse clinical contexts, comparison of additional LLMs, and formal assessment of prior informativeness and clinical impact. Integration with regulatory frameworks and extension to other types of clinical endpoints are also important directions for further work.
@@ 6 Conclusion and Future Work
This study demonstrates that LLM-informed prior elicitation, using models such as Llama 3.3 and MedGemma, can substantially improve predictive performance and sample efficiency in hierarchical Bayesian modeling for clinical trials. Through rigorous cross-validation and sample efficiency experiments on real HRPC trial data, we showed that LLM-derived priors consistently outperform traditional meta-analytical baselines, even when the available training data is limited. Llama 3.3 (Blind, $T=1.0$) provided the best overall results, and both LLMs enabled robust inference with fewer patients, highlighting the practical value of generative AI for clinical safety analysis.
@@ Appendix E Full Disease-Informed Prompt
3TASK: Provide ONLY rate parameters for exponential priors based on HRPC control arm data.
@@
6- Disease: Metastatic hormone-resistant prostate cancer (HRPC)
@@
17- Use your expert knowledge and draw on published clinical trials, empirical data, or established domain knowledge specific to HRPC control arms to set informative (not weakly-informative or non-informative) prior rates.
18- Avoid using vague or default values. Base your answer on realistic HRPC control arm data or strong prior experience relevant to typical AE rates in multi-center oncology trials.